# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 8.2 BRADSO Incentive Program and the USMA-2006 Mechanism
Switzer’s research strategy in his thesis, which followed the basic structure of Balinski and Sönmez (1999), made sense because there are many parallels between the U.S. Army’s branching process and the Turkish student placement (cf. Sections 6.1-6.6). The Army’s branching process involves assigning USMA cadets to military specialties—branches—in their senior year. Branching among the 17 specialties (e.g., Aviation, Infantry, Military Intelligence, etc.) plays an important role in a cadet’s career progression.
@@
Under SSD-OML, cadets submit their preference rankings for branches, and the highest OML-ranked cadet receives their first-choice branch. The next highest OML-ranked cadet receives their top choice among the remaining positions, and so on. Therefore, cadet claims over positions at Army branches depended solely on their performance rankings before 2006—a practice that makes sense given the importance of hierarchy in the Army. Moreover, due to the importance of trust among cadets themselves and between cadets and the USMA leadership, another key advantage of this mechanism was its strategy-proofness. Naturally, a system that allows cadets to be outmaneuvered by others or by the system itself would not foster trust among its participants.
To address historically low retention rates among junior officers, the Army introduced a second pathway for cadets to access their preferred branches, starting with the Class of 2006. This initiative allowed cadets to receive elevated priority for a portion of positions in any branch they selected, provided they committed to an additional three years of Active Duty Service Obligation (ADSO) beyond the standard requirement. At USMA, the standard ADSO is five years. This incentive program became known as the Branch-of-Choice Active Duty Service Obligation, or simply the BRADSO program.
Given the opposition from many senior Army officers to interfering with a system that allocates positions solely based on performance ratings, this second pathway to branch access was limited, as a compromise, to 25% of the positions in each branch at USMA. [^92] See, for example, the following quote from Appendix E of Colarusso et al. (2010): “The branch and post incentives also raised concerns. Devoted supporters of the ROTC and West Point Order of Merit (OML) system for allocating branches and posts objected that low OML cadets could ’buy’ their branch or post of choice ahead of higher OML cadets. Since branch and post assignments represent a zero-sum game, the ability of cadets with a lower OML ranking to displace those above them was viewed by some as unfair or as undermining the OML system. However, rather than undermining the legacy system or creating inequities, the branch and post incentives program makes willingness to serve a measure of merit in branching and posting, thus providing taxpayers a fair return on their officer accessions investment.” Priority for the remaining 75% of the positions at each branch continued to be based on the OML.
@@
Under the USMA-2006 mechanism, cadets were asked to rank their preferences over branches and select a set of branches for which they would invoke the BRADSO program to receive preferential access to the last 25% of positions. Based on the submitted strategies of cadets, a procedure similar to SSD-OML was used to allocate positions. However, unlike this procedure, once the initial 75% of positions in a given branch were assigned, the second criterion was applied to allocate the remaining positions. A cadet who received a position in a given branch, say Infantry, was charged an additional three years of ADSO only if (i) they were assigned one of the last 25% of Infantry positions and (ii) had volunteered for the Infantry BRADSO incentive.
As natural as this modification sounds, it has several failures that were discovered in Switzer (2011). Most notably, cadets could receive less-preferred outcomes than some of their peers despite having higher OML rankings, merely because they volunteered for the BRADSO incentive for their assigned branches. Formally, the USMA-2006 mechanism fails to satisfy the following axiom. [^93] For the special case of student placement, this axiom reduces to no justified envy. Following the terminology in Balinski and Sönmez (1999), this axiom is also referred to as fairness in Switzer (2011) and Sönmez and Switzer (2013).
@@
The failure of no priority reversal under the USMA-2006 mechanism reflects an inability to harmoniously integrate two allocation criteria in the design. For the same reason, the USMA-2006 mechanism also creates incentives to game the system, either by misreporting branch preferences or by concealing willingness to participate in the BRADSO program.
@@
Fortunately, addressing this issue is relatively straightforward with a simple adjustment to the strategy space. Instead of asking cadets to submit their preferences over branches along with the set of branches for which they volunteer for the BRADSO incentive, cadets could be asked to submit their preferences over branch-price pairs. This adjustment would resolve the issue.
@@ 8.3 Switzer’s Reform Proposal for the USMA-2006 Mechanism
To address the first root cause of failures in the USMA-2006 mechanism, Switzer (2011) modifies the strategy space by having cadets submit their preferences for branch-price pairs, as discussed.
To tackle the more complex second root cause of failures, Switzer (2011) introduces a priority ranking of cadet-price pairs for each branch, where:
@@
    Any pair with an increased price has higher priority than any pair with the base price, regardless of the cadets in those pairs.
@@
With this priority structure, the outcome function of Switzer’s proposed mechanism follows a procedure reminiscent of the SSD-OML. However, unlike SSD-OML, cadets submit their applications to branches along with their proposed prices. Branches then use Switzer’s constructed priorities to evaluate these applications. Since a cadet can submit two applications for any given branch—one at the base price and another at the increased price—a branch may reject a cadet who was tentatively admitted at the base price in favor of another cadet applying with the increased price. This introduces an iterative structure similar to the McVitie and Wilson (1970) version of Gale and Shapley’s celebrated Deferred Acceptance algorithm, where individual proposals are made sequentially rather than simultaneously.
It is clear that Switzer’s procedure terminates in a finite number of steps, and the resulting mechanism satisfies both no priority reversal and strategy-proofness, effectively addressing the shortcomings of the USMA-2006 mechanism. From a purely research standpoint, this is a well-behaved mechanism. However, from a policy perspective, Switzer’s proposal potentially disrupts some delicate balances.
Under Switzer’s proposed priority structure, all units within a given branch use the same criteria for allocation, with price always being the primary consideration. This specific design choice was the source of my skepticism regarding the mechanism’s viability for the Army. As discussed in Section 8.2, the Army’s policy maintained a delicate balance by granting BRADSO volunteers elevated priority for only a fraction of available positions under the USMA-2006 mechanism, not all. At the time, although I didn’t know the exact rationale behind this limitation, I strongly believed it wasn’t arbitrary. [^94] As evidenced by Footnote 92, this hunch turned out to be correct. This is why I viewed this aspect of Switzer’s mechanism as a potential “deal-breaker” and chose to refine it by incorporating both criteria, as the Army had done. In my view, while Switzer’s mechanism successfully addressed some critical shortcomings of the USMA-2006 mechanism, it would likely introduce new ones.
@@ 8.4 Leveraging Matching with Contracts for a Minimalist Reform
I found myself wondering: Why did Switzer not incorporate the Army’s dual-criteria priority structure into his proposed mechanism? After all, the remaining elements of his mechanism could still have been used with this modification, making it much closer to the USMA-2006 mechanism.
After reflecting on this for a while, I made several observations that put Switzer’s proposal into perspective. The version of the mechanism incorporating the Army’s dual-criteria priority structure is significantly more complex to analyze than the single-criterion version. By contrast, under Switzer’s constructed priority structure, the analysis become more manageable.
The key to advancing the “discovery-invention cycle”—which began with Switzer’s thesis and ultimately led to a reform of the Army’s branching process—was a series of breakthroughs in basic theory from the mid-2000s to the early 2010s.
The Army’s branching process, as studied in Switzer (2011), can be framed as a special case of the celebrated matching with contracts model by Hatfield and Milgrom (2005), a generalization of Gale and Shapley’s many-to-one matching model discussed earlier in Section 6.4. In their generalization, Hatfield and Milgrom allow agents (e.g., students or cadets) and institutions (e.g., colleges or branches) to be matched under multiple arrangements, introducing the concept of a contractual term between them. Each agent’s preferences are expressed over their own contracts—triples that specify the agent, an institution, and the contractual term. Each institution, in turn, evaluates its set of contracts using a choice rule, a function that selects a set of feasible contracts for the institution from any given set of contracts.
Hatfield and Milgrom (2005) assume that each institution’s choice rule satisfies the following two technical conditions. [^95] The assumption of the IRC condition in Definition 12 is implicit in Hatfield and Milgrom (2005); see Aygün and Sönmez (2013) for details.
@@
A choice rule $C(.)$ satisfies the substitutability condition if, whenever a contract is selected from a set of contracts, it remains selected from any smaller subset that results from the removal of other contracts originally available.
@@
A choice rule $C(.)$ satisfies the independence of rejected contracts (IRC) condition if the removal of an unselected contract from the set of available contracts never affects the set of selected contracts.
Under these assumptions, Hatfield and Milgrom (2005) generalized Gale and Shapley’s celebrated individual-proposing Deferred Acceptance algorithm through the following iterative algorithm they called the cumulative offer process:
At each step, and following any given priority ranking of agents—which becomes immaterial to the final outcome under the substitutability and IRC conditions—the highest-priority agent without a contract currently “on hold” proposes their most-preferred, previously unrejected contract to the institution associated with that contract. The institution then reviews all offers it has received up to that point (including those previously rejected), retains the best set of contracts based on its choice rule, and rejects the rest.
As long as all choice rules satisfy the substitutability and IRC conditions, the algorithm ensures that no previously rejected contract is held at a later step, guaranteeing termination in a finite number of steps. The outcome of the cumulative offer process is the final set of contracts held at its conclusion.
@@
A choice rule $C(.)$ satisfies the law of aggregate demand (LAD) condition if a set-wise expansion of available contracts never results in the selection of a strictly smaller number of contracts.
@@
Here is why the analysis becomes more complex under the Army’s dual-criteria priority structure in the USMA-2006 mechanism. In this setting, the resulting branch choice rules no longer satisfy the substitutability condition, placing the application beyond the scope of Hatfield and Milgrom (2005).
Consider a set of cadet-price pairs that includes both the base-price and increased-price options for a given cadet, $c$. A branch might select the increased-price option and reject the base-price option, but in a smaller, less competitive set, the branch could reverse this decision, choosing to reject the increased-price option instead.
In addition to falling outside the scope of Hatfield and Milgrom (2005), the setting with the Army’s dual-criteria priority structure also lies beyond the reach of Switzer’s proposed mechanism, as his solution relies on a uniform priority ranking of cadet-price pairs for all positions within any given branch.
When I received Switzer’s email in May 2011, I initially thought it represented a direct application of Hatfield and Milgrom (2005). This would have been an exciting development for theory, as it would mark the first real-world application of the matching with contracts model beyond its special case under Gale and Shapley’s many-to-one matching model without any contractual terms. At first glance, the cumulative offer process appeared to work seamlessly with the Army’s dual-criteria priority structure, and the resulting mechanism seemed to address the shortcomings of the USMA-2006 mechanism.
@@
Fortunately, recent research in pure theory by Hatfield and Kojima (2008, 2010) provided the much-needed clarification. Contrary to a claim in Hatfield and Milgrom (2005), Hatfield and Kojima (2008) showed that the substitutability condition is not necessary for a well-defined and well-behaved cumulative offer process. Moreover, thanks to an important generalization of Hatfield and Milgrom (2005) by Hatfield and Kojima (2010), the cumulative offer mechanism continues to satisfy both no priority reversal and strategy-proofness, as long as the choice rules satisfy the following weaker substitutability condition, along with the IRC and LAD conditions.
@@
A choice rule $C(.)$ satisfies the unilateral substitutability condition if, whenever the only available contract of a cadet is selected from a set of contracts, it remains selected from any smaller subset that results from the removal of other contracts originally available.
@@
In Sönmez and Switzer (2013), we have shown that, even though the choice rules induced by the Army’s dual-criteria priority structure do not satisfy the substitutes condition, they satisfy the milder unilateral substitutability condition. In addition, they satisfy the IRC and LAD conditions as well, putting everything in place.
@@
As a result, when the cumulative offer mechanism is implemented under the Army’s dual-criteria priority structure, it remains both well-defined and well-behaved. A modest adjustment to the mechanism’s strategy space, combined with a natural adaptation of its outcome function, successfully addresses the shortcomings of the USMA-2006 mechanism—crucially, without altering the Army’s original priority structure. Sönmez and Switzer (2013) proposes the resulting multi-price cumulative offer (MPCO) mechanism as a minimalist enhancement of the USMA-2006 mechanism.

### v4

@@ 8.2 BRADSO Incentive Program and the USMA-2006 Mechanism
Switzer’s research strategy in his thesis, which followed the basic structure of Balinski and Sönmez (1999), was sensible given the many parallels between the U.S. Army’s branching process and the Turkish student placement system (cf. Sections 6.16.6). The Army’s branching process assigns USMA cadets to military specialties—known as branches—during their senior year. Branching across the 17 specialties (e.g., Aviation, Infantry, Military Intelligence, etc.) plays a central role in shaping a cadet’s career trajectory.
@@
Under SSD-OML, cadets submit their preference rankings for branches, and the highest OML-ranked cadet receives their first-choice branch. The next highest-ranked cadet receives their top choice among the remaining positions, and so on. Thus, cadet claims over positions at Army branches depended solely on their performance rankings prior to 2006—a practice that makes sense given the importance of hierarchy in the Army. In addition, because of the importance of trust both among cadets and between cadets and the USMA leadership, another key advantage of this mechanism was its strategy-proofness. Naturally, a system that allows cadets to be outmaneuvered by others or by the system itself would not foster trust among its participants.
Starting with the Class of 2006, the Army introduced a second pathway for cadets to access their preferred branches in order to address historically low retention rates among junior officers. This initiative allowed cadets to receive elevated priority for a portion of positions in any branch they selected, provided they committed to an additional three years of Active Duty Service Obligation (ADSO) beyond the standard requirement. At USMA, the standard ADSO is five years. The incentive program became known as the Branch-of-Choice Active Duty Service Obligation, or simply the BRADSO program.
Given the opposition from many senior Army officers to interfering with a system that allocates positions solely based on performance ratings, this second pathway to branch access was limited, as a compromise, to 25% of the positions in each branch at USMA.104104104See, for example, the following quote from Appendix E of Colarusso et al. (2010): “The branch and post incentives also raised concerns. Devoted supporters of the ROTC and West Point Order of Merit (OML) system for allocating branches and posts objected that low OML cadets could ’buy’ their branch or post of choice ahead of higher OML cadets. Since branch and post assignments represent a zero-sum game, the ability of cadets with a lower OML ranking to displace those above them was viewed by some as unfair or as undermining the OML system. However, rather than undermining the legacy system or creating inequities, the branch and post incentives program makes willingness to serve a measure of merit in branching and posting, thus providing taxpayers a fair return on their officer accessions investment.” Priority for the remaining 75% of the positions at each branch continued to be based on the OML.
@@
USMA-2006 Mechanism.
Cadets rank their branch preferences and specify the branches in which they would invoke BRADSO incentives to obtain preferential access to the final 25% of positions within those branches. Based on the submitted profile of cadet strategies, a procedure similar to SSD-OML is used to allocate positions. However, unlike SSD-OML, once the initial 75% of positions in a given branch are assigned, a second criterion is applied to allocate the remaining positions. A cadet who receives a position in a given branch—say, Infantry—is charged an additional three years of ADSO only if (i) they are assigned one of the last 25% of Infantry positions and (ii) they had volunteered for the Infantry BRADSO incentive.
As natural as this modification sounds, it has several failures identified in Switzer (2011). Most notably, cadets could receive less-preferred outcomes than some of their peers despite having higher OML rankings, merely because they volunteered for the BRADSO incentive in their assigned branches. Formally, the USMA-2006 mechanism fails to satisfy the following axiom.105105105For the special case of student placement, this axiom reduces to no justified envy. Following the terminology in Balinski and Sönmez (1999), this axiom is also referred to as fairness in Switzer (2011) and Sönmez and Switzer (2013).
@@
The failure of no priority reversal under the USMA-2006 mechanism reflects an inability to properly integrate two allocation criteria into the design. For the same reason, the USMA-2006 mechanism also creates incentives to game the system, either by misreporting branch preferences or by concealing their willingness to volunteer for BRADSO incentives.
@@
Fortunately, addressing this issue is relatively straightforward with a simple adjustment to the strategy space. Instead of asking cadets to submit their preferences over branches along with the set of branches for which they volunteer for the BRADSO incentive, cadets could be asked to submit their preferences over branch-price pairs. This minimalist adjustment would resolve the issue.
@@ 8.3 Switzer’s Reform Proposal for the USMA-2006 Mechanism
To address the first root cause of failures in the USMA-2006 mechanism, Switzer (2011) modifies the strategy space as discussed above so that cadets submit their preferences for branch–price pairs.
To tackle the more complex second root cause, Switzer (2011) introduces a priority ranking of cadet–price pairs for each branch:
@@
    Any pair with an increased price has higher priority than any pair with the base price, regardless of which cadets are involved.
@@
Under this priority structure, the outcome function of Switzer’s proposed mechanism follows a procedure reminiscent of the SSD-OML. However, unlike SSD-OML, cadets submit their applications to branches along with their proposed prices. Branches then apply Switzer’s priority structure to evaluate these applications. Because a cadet may rank the same branch at both the base and the increased price, a branch may reject a cadet tentatively admitted at the base price in favor of another cadet applying at the increased price. This introduces an iterative structure similar to the McVitie and Wilson (1970) variant of the deferred acceptance algorithm, in which proposals proceed sequentially rather than simultaneously as in Gale and Shapley’s original version.
It is clear that Switzer’s procedure terminates in a finite number of steps, and the resulting mechanism satisfies both no priority reversal and strategy-proofness, effectively addressing the shortcomings of the USMA-2006 mechanism. From a research standpoint, his design is well-behaved. However, from a policy perspective, the proposal may disrupt some delicate balances.
Under Switzer’s proposed priority structure, all units within a given branch use the same criteria for allocation, with price always being the primary consideration. This design choice was the source of my skepticism regarding its viability for the Army. As discussed in Section 8.2, the Army’s policy maintained a delicate balance by granting BRADSO volunteers elevated priority for only a fraction of available positions under the USMA-2006 system, not all. Although I didn’t know the exact rationale at the time, I strongly believed the limitation wasn’t arbitrary.106106106As evidenced by Footnote 104, this hunch turned out to be correct. For this reason, I viewed Switzer’s design as a potential “deal-breaker” and chose to refine it by incorporating both criteria, as the Army had done. In my view, while Switzer’s proposal successfully addressed some critical shortcomings of the USMA-2006 mechanism, it would likely introduce new ones.
@@ 8.4 Leveraging Matching with Contracts for a Minimalist Reform
I found myself wondering why Switzer did not incorporate the Army’s dual-criteria priority structure into his proposed mechanism. After all, the remaining elements of his proposal could still have been used with this modification, making it more closely aligned with the USMA-2006 mechanism.
After some reflection, I made a number of observations that put Switzer’s proposal into perspective. Incorporating the Army’s dual-criteria priority structure makes the mechanism significantly more complex to analyze than the single-criterion version. By contrast, under Switzer’s constructed priority structure, the analysis becomes more manageable.
As illustrated in Figure 36, the key to advancing the ongoing “discovery–invention cycle”—which expanded through Switzer’s thesis and ultimately culminated in a reform of the Army’s branching process—was a series of breakthroughs in basic theory from the mid-2000s to the early 2010s.
The Army’s branching process, as studied in Switzer (2011), can be framed as a special case of the celebrated matching with contracts model by Hatfield and Milgrom (2005), a generalization of Gale and Shapley’s many-to-one matching model discussed earlier in Section 6.4. In their generalization, Hatfield and Milgrom allow agents (e.g., students or cadets) and institutions (e.g., colleges or branches) to be matched under multiple arrangements, introducing the concept of a contractual term between them. Each agent’s preferences are expressed as preferences over contracts—triples that specify the agent, an institution, and the contractual term. Each institution, in turn, evaluates its set of contracts using a choice rule, a function that selects a set of feasible contracts for the institution from any given set of contracts.
Hatfield and Milgrom (2005) assume that each institution’s choice rule satisfies the following two technical conditions.107107107The assumption of the IRC condition in Definition 12 is implicit in Hatfield and Milgrom (2005); see Aygün and Sönmez (2013) for details.
@@
A choice rule $C(\cdot)$ satisfies substitutability if, whenever a contract is chosen from a set of contracts, it continues to be chosen from every subset obtained by removing other contracts from that set.
@@
A choice rule $C(\cdot)$ satisfies the independence of rejected contracts (IRC) condition if removing an unselected contract from the set of available contracts does not change the set of selected contracts.
Under these assumptions, Hatfield and Milgrom (2005) generalized Gale and Shapley’s celebrated individual-proposing deferred acceptance algorithm through the following iterative algorithm.
Cumulative Offer Process.
At each step, following any given priority ranking of agents—which is immaterial to the final outcome under the substitutability and IRC conditions—the highest-priority agent without a contract “on hold” proposes their most-preferred contract that has not yet been rejected to the institution associated with that contract. The institution then reviews all offers it has received up to that point, including those it previously rejected, retains the best set of contracts based on its choice rule, and rejects the rest. The process ends when no new proposal is made.
As long as all choice rules satisfy the substitutability and IRC conditions, the algorithm ensures that no previously rejected contract is ever held at a later step, thereby guaranteeing termination in a finite number of steps.
@@
A choice rule $C(\cdot)$ satisfies the law of aggregate demand (LAD) if, whenever the set of available contracts expands, the number of contracts selected weakly increases.
@@
The analysis becomes more complex under the Army’s dual-criteria priority structure in the USMA-2006 mechanism because the resulting branch choice rules no longer satisfy substitutability, placing the application beyond the technical scope of Hatfield and Milgrom (2005).
To illustrate, consider a set of contracts that includes both the base-price and increased-price options for a cadet $c$. A branch might select the increased-price contract and reject the base-price one, but in a smaller set with less competition it could reverse this decision, instead choosing the base-price contract and rejecting the increased-price one.
In addition to falling outside the scope of Hatfield and Milgrom (2005), the Army’s dual-criteria priority structure also lies beyond the reach of Switzer’s proposed mechanism, since his solution relies on a uniform priority ranking of cadet–price pairs for all positions within a branch.
When I received Switzer’s email in May 2011, I initially thought the Army’s branching process was a direct application of Hatfield and Milgrom (2005). This would have been an exciting development for theory, as it would mark the first real-world application of the matching-with-contracts model, beyond its special case of Gale and Shapley’s many-to-one matching without contractual terms. At first glance, the cumulative offer process appeared to work seamlessly with the Army’s dual-criteria priority structure, and the resulting mechanism seemed to address the shortcomings of the USMA-2006 mechanism.
@@
Fortunately, recent research in pure theory by Hatfield and Kojima (2008, 2010) provided the much-needed clarification. Contrary to a claim in Hatfield and Milgrom (2005), Hatfield and Kojima (2008) showed that the substitutability condition is not necessary for a well-defined and well-behaved cumulative offer process. Moreover, thanks to the important generalization of Hatfield and Milgrom (2005) developed by Hatfield and Kojima (2010), the cumulative offer mechanism continues to satisfy both no priority reversal and strategy-proofness, as long as the choice rules satisfy the following weaker substitutability condition, along with the IRC and LAD conditions.
@@
A choice rule $C(\cdot)$ satisfies the unilateral substitutability condition if, whenever the only available contract of a cadet is chosen from a set of contracts, it continues to be chosen from every subset obtained by removing other contracts from that set.
@@
Although the choice rules induced by the Army’s dual-criteria priority structure fail to satisfy the substitutes condition, they satisfy the milder unilateral substitutability condition as well as IRC and LAD, thus ensuring that the cumulative offer process is well-defined and well-behaved for the Army’s branching process.
Multi-Price Cumulative Offer (MPCO) Mechanism.
@@
Building on this foundation, and adopting the cumulative offer process under the Army’s dual-criteria priority structure, the MPCO mechanism modifies USMA-2006 by introducing a modest adjustment to the strategy space and a natural adaptation of the outcome function. These changes resolve the shortcomings of USMA-2006—crucially, without altering the Army’s original priority structure. As a minimalist enhancement of the USMA-2006 mechanism, Sönmez and Switzer (2013) proposes the MPCO for the Army.108108108Alternatively, the MPCO mechanism can be constructed as a descending salary adjustment process in the Kelso and Crawford (1982) model—a key precursor to Hatfield and Milgrom (2005)—as shown by Jagadeesan (2019).
We next execute the MPCO mechanism and the cumulative offer process using an example, as depicted in Figure 22.
Example 9.
There are five cadets—Alp, Banu, Cora, Diya, and Ezra—with earlier names in the alphabet corresponding to higher rankings in the Order of Merit List (OML). There are three branches: Infantry ($I$), Military Intelligence ($M$), and the Quartermaster Corps ($Q$), with two positions each in Infantry and Military Intelligence, and one in the Quartermaster Corps.
The first positions in Infantry and Military Intelligence—denoted as $I_{r}$ and $M_{r}$—and the only position in the Quartermaster Corps are offered only at the base level of Active Duty Service Obligation (ADSO) (also called base price) and the priorities for these positions are determined solely by the OML. These are called regular positions.
The second positions in the Infantry and Military Intelligence—denoted as $I_{f}$ and $M_{f}$—are called flexible ADSO (or flexible-price) positions. As part of the BRADSO program, these flexible-price positions give priority access to cadets who indicate a willingness to extend their ADSO by three years and rely on the OML among cadets with the same ADSO willingness. These flexible-price positions may be awarded with either the base or the extended ADSO.
For the Quartermaster Corps, the choice rule is based solely on the OML, admitting the highest-ranking applicant. For branches with both regular and flexible-price positions—Infantry and Military Intelligence—the choice rules rely on the slot-specific priorities indicated above, filling first the regular position and then the flexible-price position.
For any cadet and branch, let the superscript 0 indicate the contract with the base ADSO, and the superscript + indicate the contract with the extended ADSO. For example, a contract between Alp and Infantry at the base ADSO is written as $I^{0}$ in Alp’s preferences (or when he proposes it), and as A0 in the priority ordering of an Infantry position (or when Infantry receives it as a proposal).
Using this notation, cadet preferences and the slot-specific priorities of the positions over acceptable contracts are given as follows:
$\begin{array}[]{llll}\mbox{A}:&I^{0}-M^{0}-Q^{0}\hskip 18.49988pt\qquad&I_{r}:&\mbox{A}^{0}-\mbox{B}^{0}-\mbox{C}^{0}-\mbox{D}^{0}-\mbox{E}^{0}\\ \mbox{B}:&I^{0}-I^{+}-Q^{0}-M^{0}&I_{f}:&\mbox{A}^{+}-\mbox{B}^{+}-\mbox{C}^{+}-\mbox{D}^{+}-\mbox{E}^{+}-\mbox{A}^{0}-\mbox{B}^{0}-\mbox{C}^{0}-\mbox{D}^{0}-\mbox{E}^{0}\\ \mbox{C}:&I^{0}-M^{0}-I^{+}-Q^{0}\hskip 18.49988pt\qquad&M_{r}:&\mbox{A}^{0}-\mbox{B}^{0}-\mbox{C}^{0}-\mbox{D}^{0}-\mbox{E}^{0}\\ \mbox{D}:&I^{0}-I^{+}-Q^{0}-M^{0}&M_{f}:&\mbox{A}^{+}-\mbox{B}^{+}-\mbox{C}^{+}-\mbox{D}^{+}-\mbox{E}^{+}-\mbox{A}^{0}-\mbox{B}^{0}-\mbox{C}^{0}-\mbox{D}^{0}-\mbox{E}^{0}\\ \mbox{E}:&M^{0}-I^{0}-M^{+}-Q^{0}\hskip 18.49988pt\qquad&Q:&\mbox{A}^{0}-\mbox{B}^{0}-\mbox{C}^{0}-\mbox{D}^{0}-\mbox{E}^{0}\end{array}$
We now process the cumulative offer process, at each step considering the highest-OML cadet without a contract held by any branch.
Step 1. Alp proposes to Infantry his base ADSO contract A0. Infantry starts its evaluation with its regular position $I_{r}$ and holds the offer for that position.
Step 2. Banu proposes to Infantry her base ADSO contract B0. Infantry starts its evaluation with its regular position $I_{r}$, considers it together with A0, and rejects it for this position, continuing to hold A0. It then evaluates B0 for its flexible-price position $I_{f}$ and holds it for that position.
Step 3. Cora proposes to Infantry her base ADSO contract C0. Infantry starts its evaluation with its regular position $I_{r}$, considers it together with its earlier cumulative offers A0 and B0, and rejects it for this position along with B0, holding A0. It then evaluates C0 for its flexible-price position $I_{f}$ together with B0 and rejects it, continuing to hold B0.
Step 4. Cora proposes to Military Intelligence her base ADSO contract C0. Military Intelligence starts its evaluation with its regular position $M_{r}$ and holds the offer for that position.
Step 5. Diya proposes to Infantry her base ADSO contract D0. Infantry starts its evaluation with its regular position $I_{r}$, considers it together with its earlier cumulative offers A0, B0, and C0, and rejects it for this position along with B0 and C0, holding A0. It then evaluates D0 for its flexible-price position $I_{f}$ together with B0 and C0 and rejects it along with C0, continuing to hold B0.
Step 6. Having just been rejected by Infantry with her base ADSO contract, Diya next proposes to Infantry her extended ADSO contract D+. Ineligible for the regular position $I_{r}$, Infantry evaluates it for the flexible-price position $I_{f}$ together with its earlier cumulative offers B0, C0, and D0, holding D+ and rejecting the others, including B0, which it had been holding since Step 2.
Step 7. Having been rejected by Infantry with her base ADSO contract after being on hold for several steps, Banu next proposes to Infantry her extended ADSO contract B+. Ineligible for the regular position $I_{r}$, Infantry evaluates it for the flexible-price position $I_{f}$ together with its earlier cumulative offers B0, C0, D0, and D+, holding B+ and rejecting the others, including D+, which it had been holding in the previous step.
Step 8. Having been rejected by Infantry with her extended ADSO contract as well, Diya next proposes to the Quartermaster Corps her base ADSO contract D0. The Quartermaster Corps holds it for its only position.
Step 9. Ezra proposes to Military Intelligence his base ADSO contract E0. Military Intelligence starts its evaluation with its regular position $M_{r}$, considers it together with C0, and rejects E0, continuing to hold C0. It then evaluates E0 for its flexible-price position $M_{f}$ and, having no offer held there, holds E0.
With no cadet rejected in Step 9, the process ends. The contracts on hold are finalized: Alp receives the regular position at Infantry with the base ADSO, Banu receives the flexible-price position at Infantry with the extended ADSO, Cora receives the regular position at Military Intelligence with the base ADSO, Diya receives the position at Quartermaster Corps with the base ADSO, and Ezra receives the flexible-price position at Military Intelligence with the base ADSO.
Figure 22: Execution of the MPCO mechanism in Example 9. At each step, the highest OML-ranked applicant makes a proposal to their highest-ranked un-rejected contract. Offers are shown in boxes, while earlier proposals that have been rejected are shown in faint gray and struck out with a faint gray line. For each branch that receives a new proposal, cumulative offers to each slot remain boxed, with rejected ones struck out.